OpenStudy (anonymous):

5^-2x=3 it says solve for x and use natural log where ever log is needed the answer is -ln3/2ln5 I have all the log properties in front of me and still cant make sense of this stuff!

OpenStudy (anonymous):

take the ln of both sides you should get -2x*ln(5)=ln(3) This is because when you take log or ln, you can move the exponents down to the front and multiply them so solve equation: -2x = ln3/ln5 x= ln3/(-2*ln5)

OpenStudy (anonymous):

ok thanks alot!